OpenStudy (anonymous):

If a radioactive nuclide has a half-life of 1.3 billion years and it is determined that a rock contains 25% of its original amount of this nuclide, about how old is the rock? 1.3 billion years 2.6 billion years 0.65 billion years 3.9 billion years

OpenStudy (anonymous):

\[\large y=Ce^{kt}\] where \(y\) is the amount of the radioactive material at time \(t\), \(C\) is the initial amount (at \(t=0\)), \(k\) is the relative decay factor, and \(t\) is time (in billions of years). We know the half-life is 1.3 billion years, which means after 1.3 billion years, we have half of the initial amount: \[\large y=\frac{C}{2}=Ce^{1.3k}~~\Rightarrow~~\frac{1}{2}=e^{1.3k}\] Solve for \(k\): \[\large\begin{align*} \frac{1}{2}&=e^{1.3k}\\ \ln\frac{1}{2}&=1.3k\\ k&=\frac{1}{1.3}\ln\frac{1}{2}\\ k&\approx -0.533 \end{align*}\] We want to find out how old the rock is that contains 25% of its original amount. In other words, solve for \(t\): \[\large 0.25=e^{kt}\] where \(k\approx-0.533\).